Please help me with this question: "M is the midpoint of YZ. If YM=x+3 and YZ=3x-1, find MZ.
We know ym = 1/2(yz) Plug in teh numbers we know
x + 3 = 1/2(3x - 1) Solve
x + 3 = 1.5x - .5 Subtract 1.5x on both sides and subtract 3.
-.5x = -3.5 Divide -.5 on both sides
x = 7 Plug 7 into x + 3 and you have 10 = YM since It is the midpoint then MZ is also 10
RJ Toftness
www.math-unlock.com

Found 3 solutions by user_dude2008, unlockmath, stanbon:
Answer by user_dude2008(1825)

YM=MZ ----> YM+MZ=YZ

YM+MZ=YZ
(x+3)+MZ=3x-1
MZ=3x-1-x-3
MZ=2x-4

YM=MZ
x+3=2x-4
3+4=2x-x
7=x
x=7

MZ=2(7)-4=14-4=10

Ans: MZ=10

Answer by unlockmath(1122)

We know ym = 1/2(yz) Plug in teh numbers we know
x + 3 = 1/2(3x - 1) Solve
x + 3 = 1.5x - .5 Subtract 1.5x on both sides and subtract 3.
-.5x = -3.5 Divide -.5 on both sides
x = 7 Plug 7 into x + 3 and you have 10 = YM since It is the midpoint then MZ is also 10
RJ Toftness
www.math-unlock.com

Answer by stanbon(48558)

Please help me with this question: "M is the midpoint of YZ. If YM=x+3 and YZ=3x-1, find MZ.
-----------------
Draw the picture:
If M is the midpoint, YM + MZ = yZ
Substitute and solve for MZ:
(x+3)+MZ = 3x-1
MZ = (3x-1)-(x+3)
MZ = 2x - 4
=================
But M is the midpoint, so yM = MZ
x+3 = 2x-4
x = 7
---
If x = 7 then MZ = 2x-4 = 2*7-4 = 10
========================================
Cheers,
Stan H.
Cheers,
Stan H.